OpenStudy (anonymous):

Katrina is fencing her yard, which is shaped as a rhombus. She used (5x + 1) meters of fencing material for one side and (7x - 11) meters for the adjacent side. What is the perimeter of the yard? 124 meters 116 meters 140 meters 168 meters

OpenStudy (anonymous):

First of all, a rhombus has equal sides. Can you please check the question--it doesn't make sense.

Parth (parthkohli):

Not really, Rajath :) \(5x + 1 = 7x - 11\) Solve for \(x\), then find the perimeter.

OpenStudy (anonymous):

ahh, yes. Sorry.

OpenStudy (anonymous):

i think it means (5x+1)=(7x-11)

Parth (parthkohli):

First, solve for \(x\). Second, try to find \(5x + 1\)(or \(7x - 11\) according to your wish).

Parth (parthkohli):

As a rhombus has all its sides equal, and the number of sides is 4, we may just add the length of each sides 4 times. Remember that perimeter is just the sum of outer sides ;)

OpenStudy (anonymous):

x=6?

Parth (parthkohli):

Yes :)

Parth (parthkohli):

Now what is \(5(6) + 1\)?

OpenStudy (anonymous):

31

Parth (parthkohli):

Correct - that is the length of each side.

OpenStudy (anonymous):

31*4+124

Parth (parthkohli):

I think you meant 31 * 4 = 124 but it's all fine :)

OpenStudy (anonymous):

sory + means =

Parth (parthkohli):

and that is the answer!

OpenStudy (anonymous):

thanks a\[\infty \]

Parth (parthkohli):

Hahaha!

Parth (parthkohli):

By the way, you did \(31.841\pi\%\) work, so no need to thank me:)

OpenStudy (anonymous):

as long as the work you do is 68.159pi%

Parth (parthkohli):

lol